# Large Language Models for Cyber Security: A Systematic Literature Review

### v3

@@ 3.3. Application of LLMs in Information and Content Security
Phishing and scam detection. Network deception is a deliberate act of introducing false or misleading content into a network system, threatening the personal privacy and property security of users. Emails, short message service (SMS), and web advertisements are leveraged by attackers to entice users and steer them towards phishing sites, enticing them to click on malicious links (Tang et al., 2022). LLMs can generate deceptive or false information on a large scale under specific prompts (Ranade et al., 2021), making them useful for automated phishing email generation(Heiding et al., 2023; Roy et al., 2024), but compared to manual design methods, phishing emails generated by LLMs have lower click-through rates (Heiding et al., 2023). LLMs can achieve phishing email detection through prompts based on website information (Koide et al., 2024) or fine-tuning for specific email features (Roy et al., 2024; Moore et al., 2022). Spam emails often contain a large number of phishing emails. Labonne et al.’s research (Labonne and Moran, 2023) has demonstrated the effectiveness of LLMs in spam email detection, showing significant advantages over traditional machine learning methods. An interesting study (Cambiaso and Caviglione, 2023) suggests that LLMs can mimic real human interactions with scammers in an automated and meaningless manner, thereby wasting scammers’ time and resources and alleviating the nuisance of scam emails.
@@
Harmful contents detection. Social media platforms frequently face criticism for amplifying political polarization and deteriorating public discourse. Users often contribute harmful content that reflects their political beliefs, thereby intensifying contentious and toxic discussions or participating in harmful behavior (Wojcieszak et al., 2022). The application of LLMs in detecting harmful content can be divided into three aspects: detection of extreme political stances (Hanley and Durumeric, 2023; Mets et al., 2023), tracking of criminal activity discourse (Hu et al., 2023b), and identification of social media bots (Cai et al., 2024). LLMs tend to express attitudes consistent with the values encoded in the programming when faced with political discourse, indicating the complexity and limitations of LLMs in handling social topics (Hartvigsen et al., 2022). Hartvigsen et al. (Martin, 2023) generated a large-scale dataset of harmful and benign discourse targeting 13 minority groups using LLMs. Through validation, it was found that human annotators struggled to distinguish between LLM-generated and human-written discourse, advancing efforts in filtering and combating harmful contents.
@@
Steganography. Steganography, as discussed in Anderson’s work (Anderson and Petitcolas, 1998), focuses on embedding confidential data within ordinary information carriers without alerting third parties, thereby safeguarding the secrecy and security of the concealed information. Wang et al. (Wang et al., 2023b) introduced a method for language steganalysis using LLMs based on few-shot learning principles, aiming to overcome the limited availability of labeled data by incorporating a small set of labeled samples along with auxiliary unlabeled samples to improve the efficiency of language steganalysis. This approach significantly improves the detection capability of existing methods in scenarios with few samples. Bauer et al. (Bauer et al., 2022) used the GPT-2 model to encode ciphertext into natural language cover texts, allowing users to control the observable format of the ciphertext for covert information transmission on public platforms.
@@ 3.4. Application of LLMs in Hardware Security
Hardware vulnerability repair. LLMs have found application within the integrated System-on-Chip (SoC) security verification paradigm, showcasing potential in addressing diverse hardware-level security tasks such as vulnerability insertion, security assessment, verification, and the development of mitigation strategies (Saha et al., 2023). By leveraging hardware vulnerability information, LLMs offer advice on vulnerability repair strategies, thereby improving the efficiency and accuracy of hardware vulnerability analysis and mitigation efforts (Lin et al., 2023). In their study, Nair and colleagues (Nair et al., 2023b) demonstrated that LLMs can generate hardware-level security vulnerabilities during hardware code generation and explored their utility in generating secure hardware code. They successfully produced secure hardware code for 10 Common Weakness Enumerations (CWEs) at the hardware design level. Additionally, Tan et al. (Ahmad et al., 2023b) curated a comprehensive corpus of hardware security vulnerabilities and evaluated the performance of LLMs in automating the repair of hardware vulnerabilities based on this corpus.
@@ 3.5. Application of LLMs in Blockchain Security
Smart contract security. With the advancement of blockchain technology, smart contracts have emerged as a pivotal element in blockchain applications (Zheng et al., 2020). Despite their significance, the development of smart contracts can introduce vulnerabilities that pose potential risks such as financial losses. While LLMs offer automation for detecting vulnerabilities in smart contracts, the detection outcomes often exhibit a high rate of false positives (David et al., 2023; Chen et al., 2023d). Performance varies across different vulnerability types and is constrained by the contextual length of LLMs (Chen et al., 2023d). GPTLENS (Hu et al., 2023a) divides the detection process of smart contract vulnerabilities into two phases: generation and discrimination. During the generation phase, diverse vulnerability responses are generated, and in the discrimination phase, these responses are evaluated and ranked to mitigate false positives. Sun and colleagues (Sun et al., 2023b) integrated LLMs and program analysis to identify logical vulnerabilities in smart contracts, breaking down logical vulnerability categories into scenarios and attributes. They utilized LLMs to match potential vulnerabilities and further integrated static confirmation to validate the findings of LLMs.
@@
RQ1 - Summary (1) We have divided cybersecurity tasks into six domains: software and system security, network security, information and content security, hardware security, and blockchain security. We have summarized the specific applications of LLMs in these domains. (2) We discussed 21 cybersecurity tasks and found that LLMs are most widely applied in the field of software and system security, with 76 papers covering 8 tasks. Only 5 papers mentioned the least applied domain—blockchain security.
@@ 4.1. Architecture of LLMs in Cybersecurity
Pre-trained Language Models (PLMs) have exhibited impressive capabilities across various NLP tasks (Shanahan, 2023; Kojima et al., 2023; Minaee et al., 2024; Wei et al., 2023a; Zhao et al., 2023). Researchers have noted substantial improvements in their performance as model size increases, with surpassing certain parameter thresholds leading to significant performance gains  (Shanahan, 2023; Hoffmann et al., 2022). The term ”Large Language Model” (LLM) distinguishes language models based on the size of their parameters, specifically referring to large-sized PLMs (Minaee et al., 2024; Zhao et al., 2023).However, there is no formal consensus in the academic community regarding the minimum parameter size for LLMs, as model capacity is intricately linked to training data size and overall computational resources (Kaplan et al., 2020). In this study, we adopt to the LLM categorization framework introduced by Panet et al. (Pan et al., 2024), which classifies the predominant LLMs explored in our research into three architectural categories: encoder-only, encoder-decoder, and decoder-only. We also considered whether the related models are open-source. Open-source models offer higher flexibility and can acquire new knowledge through fine-tuning on specific tasks based on pre-trained models, while closed-source models can be directly called via APIs, reducing hardware expenses. This taxonomy and relevant models are shown in Table 5. We analyzed the distribution of different LLM architectures applied in various cybersecurity domains, as shown in Fig 5.
@@
Encoder-only LLMs. Encoder-only models, as their name implies, comprise solely an encoder network. Initially designed for language understanding tasks like text classification, these models, such as BERT and its variants (Aghaei et al., 2023b; He et al., 2021; Devlin et al., 2019; Liu et al., 2019; Ma et al., 2020; Feng et al., 2020; Guo et al., 2021; Sanh et al., 2020), aim to predict a class label for input text (Devlin et al., 2019). For instance, BERT, which adopts the encoder architecture of the Transformer model, is mentioned in 35 papers included in this study. Encoder-only LLMs use a bidirectional multi-layer self-attention mechanism to calculate the relevance of each token with all other tokens, thereby capturing semantic features that include the global context. This architecture is mainly used for processing input data, focusing on understanding and encoding information rather than generating new text. Researchers employed these models to generate embeddings for data that is relavent to cybersecurity (such as traffic data and code), mapping complex data types into vector space. These models typically use a masking strategy during pre-training, and the complex training strategies increase training time and the risk of overfitting. In the realm of cybersecurity, researchers have adopted advanced models that offer capabilities much needed in cybersecurity tasks such as code understanding (Wang et al., 2023c) and traffic analysis (Aghaei and Al-Shaer, 2023).
@@
Various prominent models, including CodeBERT (Feng et al., 2020), GraphCodeBERT (Guo et al., 2021), RoBERTa (Liu et al., 2019), CharBERT (Ma et al., 2020), DeBERTa (He et al., 2021), and DistilBERT (Sanh et al., 2020), have gained widespread usage due to their ability to effectively process and analyze code, making them valuable tools in the field of cybersecurity. An example is RoBERTa (Liu et al., 2019), which enhances BERT’s robustness through various model design adjustments and training techniques. These include altering key hyperparameters, eliminating the next-sentence pre-training objective, and utilizing substantially larger mini-batches and learning rates during training. CodeBERT (Feng et al., 2020) is a bimodal extension of BERT that utilizes both natural language and source code as its input. It employs a replaced token detection task to bolster its understanding of programming languages, in order to tackle code generation and vulnerability detection tasks. The encoder-only architecture provides models with excellent data representation capabilities. Note that these aforementioned BERT variants were not initially designed for cybersecurity tasks. Instead, their application in the cybersecurity field stems from their capabilities as general models in NLP tasks for code semantics interpretation and understanding. In contrast, SecureBERT (Aghaei et al., 2023b) is a BERT variant specifically designed for cyber threat analysis tasks. Its development highlights the robustness and flexibility of encoder-only architecture models across different tasks. Diverse training tasks and specialized training schemes enhance the model’s feature representation capabilities and boosts its performance in cybersecurity-related tasks.
@@
Regarding the model applicability, as shown in the Figure 5, encoder-only models initially garnered attention in the fields of network cybersecurity (Alkhatib et al., 2022) and software and systems cybersecurity (Lee et al., 2023; Xia and Zhang, 2022). In 2023, this concept was extended to the field of information and content cybersecurity, utilizing encoder-only models to harmful content on social media platforms (Mets et al., 2023; Hanley and Durumeric, 2023; Cai et al., 2024).
@@
Encoder-decoder LLMs. The Transformer model, based on the encoder-decoder architecture (Vaswani et al., 2023), consists of two sets of Transformer blocks: the encoder and decoder. Stacked multi-head self-attention layers are used by the encoder to encode the input sequence, generating latent representations. In contrast, the decoder performs cross-attention on these representations and sequentially produces the target sequence. The structure of encoder-decoder LLMs makes them highly suitable for sequence-to-sequence tasks such as code translation and summarization. However, their complex architecture requires more computational resources and high-quality labeled data.
Models like BART (Lewis et al., 2019), T5 (Raffel et al., 2023), and CodeT5 (Wang et al., 2021) exemplify this architecture. CodeT5 (Wang et al., 2021) and PLBART (Ahmad et al., 2021) have built upon the foundation of their original models by introducing bimodal inputs of programming language and text, demonstrating effective code comprehension capabilities. Raffle et al. (Raffel et al., 2023) show that almost all NLP tasks can be framed as a sequence-to-sequence generation task in their work. In LLM4Security, the encoder-decoder architecture was first attempted to be applied in the field of network security (Liu et al., 2020a). However, subsequent research has not widely adopted this approach, possibly due to the complexity of the encoder-decoder structure. From another perspective, owing to its flexible training strategy and excellent adaptability to complex tasks, the encoder-decoder model was later extended to other cybersecurity tasks such as program fuzzing (Deng et al., 2023c), reverse engineering (Armengol-Estapé et al., 2024), and phishing emails detection (Jamal and Wimmer, 2023).
Decoder-only LLMs. Unlike the encoder-decoder architecture, which involves the encoder processing input text and the decoder generating output text by predicting subsequent tokens from an initial state, decoder-only LLMs rely solely on the decoder module to produce the target output text (Radford et al., 2018). This autoregressive training paradigm allows decoder-only models to generate longer-form outputs token by token, making them well-suited for producing detailed analyses, advisories, and even code relevant to cybersecurity. The attention mechanism in these models also enables them to flexibly draw upon the extensive knowledge stored in their parameters and apply it to the current context.
@@
GPT-2 (Radford et al., 2019), GPT-3 (Brown et al., 2020), GPT-3.5 (OpenAI, 2022), and GPT-4 (OpenAI, 2023b) belong to the GPT series of models, among which GPT-3.5 and GPT-4 are the models most frequently used to address various cybersecurity issues in this study, covering almost all cybersecurity applications (Ferrag et al., 2024; Fayyazi and Yang, 2023; Scanlon et al., 2023; Deng et al., 2023a). Their strong few-shot learning abilities allow rapid development of new cybersecurity capabilities with minimal fine-tuning. More specialized versions like Codex (Chen et al., 2021) and others have been fine-tuned for specific code-related tasks. Open-source models like GPT-Neo (Black et al., 2021), LLaMa (Touvron et al., 2023), and Falcon (Penedo et al., 2023) also follow this architecture. Additionally, code generation LLMs such as CodeGen (Nijkamp et al., 2023), InCoder (Fried et al., 2023), StarCoder (Li et al., 2023a), and CodeLLaMa (Rozière et al., 2024) have been widely used for bug detection and repair, as well as for vulnerability repair (Wu et al., 2023a; Xia and Zhang, 2023b; Yang et al., 2023c). The large context window of decoder-only models allows them to take in and utilize more context about the cybersecurity task, like related vulnerabilities, reports, and code snippets.
@@
Due to the powerful natural language generation capabilities of the decoder-only architecture, researchers initially attempted to apply it to the generation of fake cyber threat intelligence (Ranade et al., 2021). Decoder-only LLMs have gained prominence in recent years, especially in 2022 and 2023 as shown in Figure 5, witnessing a surge in development and commercial adoption by leading Internet companies. For instance, Google introduced Bard (Google, 2023), while Meta unveiled LLaMa (Touvron et al., 2023). Unlike GPT-4 and its derivative application, ChatGPT, which quickly found integration into various cybersecurity tasks. These newer models have yet to see widespread adoption in the cybersecurity domain.

### v4

@@ 3.3. Application of LLMs in Information and Content Security
Phishing and scam detection. Network deception is a deliberate act of introducing false or misleading content into a network system, threatening the personal privacy and property security of users. Emails, short message service (SMS), and web advertisements are leveraged by attackers to entice users and steer them towards phishing sites, enticing them to click on malicious links (Tang et al., 2022). LLMs can generate deceptive or false information on a large scale under specific prompts (Ranade et al., 2021), making them useful for automated phishing email generation(Heiding et al., 2024a; Roy et al., 2024), but compared to manual design methods, phishing emails generated by LLMs have lower click-through rates (Heiding et al., 2024a). LLMs can achieve phishing email detection through prompts based on website information (Koide et al., 2024) or fine-tuning for specific email features (Roy et al., 2024; Moore et al., 2022). Spam emails often contain a large number of phishing emails. Labonne et al.’s research (Labonne and Moran, 2023) has demonstrated the effectiveness of LLMs in spam email detection, showing significant advantages over traditional machine learning methods. An interesting study (Cambiaso and Caviglione, 2023) suggests that LLMs can mimic real human interactions with scammers in an automated and meaningless manner, thereby wasting scammers’ time and resources and alleviating the nuisance of scam emails.
@@
Harmful contents detection. Social media platforms frequently face criticism for amplifying political polarization and deteriorating public discourse. Users often contribute harmful content that reflects their political beliefs, thereby intensifying contentious and toxic discussions or participating in harmful behavior (Wojcieszak et al., 2022). The application of LLMs in detecting harmful content can be divided into three aspects: detection of extreme political stances (Hanley and Durumeric, 2023; Mets et al., 2023), tracking of criminal activity discourse (Hu et al., 2024b), and identification of social media bots (Cai et al., 2024a; Thomas et al., 2024), as well as assisting human moderators in rating content against platform policies. For challenging multimodal content like memes, approaches such as EXPLAINHM leverage LLMs to generate explanatory debates between opposing personas to aid in detecting implicit harmfulness (Lin et al., 2024). LLMs tend to express attitudes consistent with the values encoded in the programming when faced with political discourse, indicating the complexity and limitations of LLMs in handling social topics (Hartvigsen et al., 2022). Hartvigsen et al. (Martin, 2023) generated a large-scale dataset of harmful and benign discourse targeting 13 minority groups using LLMs. Through validation, it was found that human annotators struggled to distinguish between LLM-generated and human-written discourse, advancing efforts in filtering and combating harmful contents.
@@
Steganography. Steganography, as discussed in Anderson’s work (Anderson and Petitcolas, 1998), focuses on embedding confidential data within ordinary information carriers without alerting third parties, thereby safeguarding the secrecy and security of the concealed information. Wang et al. (Wang et al., 2023c) introduced a method for language steganalysis using LLMs based on few-shot learning principles, aiming to overcome the limited availability of labeled data by incorporating a small set of labeled samples along with auxiliary unlabeled samples to improve the efficiency of language steganalysis. This approach significantly improves the detection capability of existing methods in scenarios with few samples. Bauer et al. (Bauer et al., 2022) used the GPT-2 model to encode ciphertext into natural language cover texts, allowing users to control the observable format of the ciphertext for covert information transmission on public platforms.
@@ 3.4. Application of LLMs in Hardware Security
Hardware vulnerability repair. LLMs have found application within the integrated System-on-Chip (SoC) security verification paradigm, showcasing potential in addressing diverse hardware-level security tasks such as vulnerability insertion, security assessment, verification, and the development of mitigation strategies (Saha et al., 2024). By leveraging hardware vulnerability information, LLMs offer advice on vulnerability repair strategies, thereby improving the efficiency and accuracy of hardware vulnerability analysis and mitigation efforts (Lin et al., 2023). In their study, Nair and colleagues (Nair et al., 2023b) demonstrated that LLMs can generate hardware-level security vulnerabilities during hardware code generation and explored their utility in generating secure hardware code. They successfully produced secure hardware code for 10 CWEs at the hardware design level. Additionally, Tan et al. (Ahmad et al., 2024a) curated a comprehensive corpus of hardware security vulnerabilities and evaluated the performance of LLMs in automating the repair of hardware vulnerabilities based on this corpus. Further investigation into automated hardware bug repair using LLMs has shown that effective prompt engineering, combined with rigorous functional and formal validation, can successfully fix security bugs in hardware description language code, outperforming synthesis-based tools (Ahmad et al., 2024b).
Hardware IP protection. Beyond detecting and repairing vulnerabilities, the role of LLMs in relation to hardware IP protection is also under investigation, including potential misuse. Research such as LLMPirate explores how LLMs might be leveraged for hardware IP piracy, demonstrating automated frameworks that employ techniques like syntax translation and interactive feedback to rewrite hardware netlists for evading detection tools, thus highlighting potential threats and the need for robust defenses (Gohil et al., 2025).
Security assertions generation. Assertion-Based Verification (ABV) using languages like SystemVerilog Assertions (SVA) is crucial for hardware security, but manually creating effective security assertions is challenging. Research is exploring the use of LLMs to automate this process; studies evaluating LLMs for generating security SVA indicate that while LLMs can produce syntactically correct assertions, the functional correctness and relevance heavily rely on the quality of the prompt context, such as providing related examples or explicit security property descriptions (Kande et al., 2024).
@@ 3.5. Application of LLMs in Blockchain Security
Smart contract vulnerability detection. With the advancement of blockchain technology, smart contracts have emerged as a pivotal element in blockchain applications (Zheng et al., 2020). Despite their significance, the development of smart contracts can introduce vulnerabilities that pose potential risks such as financial losses. While LLMs offer automation for detecting vulnerabilities in smart contracts, the detection outcomes often exhibit a high rate of false positives (David et al., 2023; Chen et al., 2024b) and require careful consideration of formal properties (Liu et al., 2025). Performance varies across different vulnerability types and is constrained by the contextual length of LLMs (Chen et al., 2024b). GPTLENS (Hu et al., 2023a) divides the detection process of smart contract vulnerabilities into two phases: generation and discrimination. During the generation phase, diverse vulnerability responses are generated, and in the discrimination phase, these responses are evaluated and ranked to mitigate false positives. Sun and colleagues (Sun et al., 2024) integrated LLMs and program analysis to identify logical vulnerabilities in smart contracts, breaking down logical vulnerability categories into scenarios and attributes. They utilized LLMs to match potential vulnerabilities and further integrated static confirmation to validate the findings of LLMs. To address risks from code reuse, SOChecker uses LLMs first to complete potentially insecure code snippets from Stack Overflow and then to identify vulnerabilities within the context of the completed code (Chen et al., 2024a). LLMs are also applied to detect specific complex semantic vulnerabilities, such as Ponzi schemes using zero-shot Chain-of-Thought prompting (Wu et al., 2024a), or accounting errors through hybrid systems combining LLM-based semantic annotation with rule-based reasoning (Zhang and Zhang, 2024). Related work also explores using LLMs to generate adversarial contracts for validating exploitability (Wu et al., 2024c) or combining LLMs with GNNs and on-chain analysis to detect deployed malicious contracts before attacks occur (Wang et al., 2024b).
@@
RQ1 - Summary (1) We have divided cybersecurity tasks into six domains: software and system security, network security, information and content security, hardware security, and blockchain security. We have summarized the specific applications of LLMs in these domains. (2)We discussed 29 cybersecurity tasks and found that LLMs are most widely applied in the field of software and system security, with 119 papers covering 11 tasks.
@@ 4.1. Architecture of LLMs in Cybersecurity
Pre-trained Language Models (PLMs) have exhibited impressive capabilities across various NLP tasks (Shanahan, 2023; Kojima et al., 2023; Minaee et al., 2024; Wei et al., 2023a; Zhao et al., 2023). Researchers have noted substantial improvements in their performance as model size increases, with surpassing certain parameter thresholds leading to significant performance gains  (Shanahan, 2023; Hoffmann et al., 2022). The term ”Large Language Model” distinguishes language models based on the size of their parameters, specifically referring to large-sized PLMs (Minaee et al., 2024; Zhao et al., 2023).However, there is no formal consensus in the academic community regarding the minimum parameter size for LLMs, as model capacity is intricately linked to training data size and overall computational resources (Kaplan et al., 2020). In this study, we adopt to the LLM categorization framework introduced by Panet et al. (Pan et al., 2024), which classifies the predominant LLMs explored in our research into three architectural categories: encoder-only, encoder-decoder, and decoder-only. We also considered whether the related models are open-source. Open-source models offer higher flexibility and can acquire new knowledge through fine-tuning on specific tasks based on pre-trained models, while closed-source models can be directly called via APIs, reducing hardware expenses. This taxonomy and relevant models are shown in Table 5. We analyzed the distribution of different LLM architectures applied in various cybersecurity domains, as shown in Fig 5.
@@
Mixtral (2) 2023.12 Yes
CodeGemma (1) 2024.09 Yes
Encoder-only LLMs. Encoder-only LLMs, exemplified by BERT (Devlin et al., 2019) (mentioned in 35 papers in this study) and its variants (Aghaei et al., 2023b; He et al., 2021; Liu et al., 2019; Ma et al., 2020; Feng et al., 2020; Guo et al., 2021; Sanh et al., 2020), primarily focus on understanding and encoding input information rather than generation. In cybersecurity, researchers employ these models primarily to generate contextualized embeddings for relevant data, such as source code, network traffic, or system logs, mapping complex data types into vector representations suitable for downstream tasks (Wang et al., 2023d; Aghaei and Al-Shaer, 2023).
@@
Various prominent models, including CodeBERT (Feng et al., 2020), GraphCodeBERT (Guo et al., 2021), RoBERTa (Liu et al., 2019), CharBERT (Ma et al., 2020), DeBERTa (He et al., 2021), and DistilBERT (Sanh et al., 2020), have gained widespread usage due to their ability to effectively process and analyze code and other security-relevant data. CodeBERT (Feng et al., 2020), for instance, is a bimodal model utilizing both natural language and source code inputs, making it suitable for tasks requiring joint understanding, such as vulnerability detection based on code and descriptions. Note that most of these aforementioned BERT variants were not initially designed for cybersecurity tasks; instead, their application in the cybersecurity field stems from their capabilities as general models adapted for code semantics interpretation and understanding. In contrast, SecureBERT (Aghaei et al., 2023b) is a BERT variant specifically designed for cyber threat analysis tasks, and other domain-specific models like HS-BERT (Meng et al., 2023) have been developed for hardware security property extraction from documentation, demonstrating the adaptation of this architecture for diverse, specialized security needs.
@@
Regarding the model applicability, as shown in the Figure 5, encoder-only models initially garnered attention in the fields of network security (Alkhatib et al., 2022) and software and systems security (Lee et al., 2023; Xia and Zhang, 2022). Specific applications include failure detection in large-scale system logs using BERT fine-tuned on normal behavior (Jeong et al., 2025), and in 2023, this concept was extended to the field of information and content security, utilizing encoder-only models like DistilBERT and RoBERTa for phishing and spam email detection (Jamal et al., 2024) and others to detect harmful content on social media platforms (Mets et al., 2023; Hanley and Durumeric, 2023; Cai et al., 2024a).
@@
Encoder-decoder LLMs. Encoder-decoder models, such as BART (Lewis et al., 2019), T5 (Raffel et al., 2023), and CodeT5 (Wang et al., 2021), structure inputs and outputs, making them suitable for sequence-to-sequence cybersecurity tasks (Vaswani et al., 2023). Specific variants like CodeT5 (Wang et al., 2021) and PLBART (Ahmad et al., 2021) utilize bimodal inputs (programming language and text), enhancing code comprehension for security analysis (Wu et al., 2023a). This architecture has demonstrated effectiveness in various security applications, including vulnerability repair, where T5-based models like VulRepair (Fu et al., 2022) achieve strong performance, although effectiveness can vary depending on the language, as noted in studies on Java vulnerability repair using models like CodeT5 and PLBART (Wu et al., 2023a). Encoder-decoder models like CodeT5 are also utilized within larger frameworks for tasks such as vulnerability detection, for instance, by generating code comments for structural analysis in SCALE (Wen et al., 2024). Furthermore, models like Flan-T5 have been specifically fine-tuned for spam and phishing email detection (Labonne and Moran, 2023; Jamal et al., 2024). Other applications include program fuzzing (Deng et al., 2024c) and assisting reverse engineering (Armengol-Estapé et al., 2024).
Decoder-only LLMs. Decoder-only LLMs utilize an autoregressive approach, generating outputs token-by-token, making them well-suited for tasks requiring detailed generation (Radford et al., 2018), such as security analyses (Jin et al., 2024), advisories (Fang et al., 2024a), and code generation (Nazzal et al., 2024).
@@
Prominent examples include the GPT series (GPT-2 (Radford et al., 2019), GPT-3 (Brown et al., 2020), GPT-3.5 (OpenAI, 2022), GPT-4 (OpenAI, 2023b)) and specialized versions like Codex (Chen et al., 2021). Among these, GPT-3.5 and GPT-4 are the models most frequently used to address various cybersecurity issues in this study, covering almost all cybersecurity applications (Ferrag et al., 2024; Fayyazi and Yang, 2023; Scanlon et al., 2023; Deng et al., 2024b). Their strong few-shot learning abilities allow rapid development of new cybersecurity capabilities with minimal fine-tuning. Open-source models like GPT-Neo (Black et al., 2021), LLaMa (Touvron et al., 2023), and Falcon (Penedo et al., 2023) also follow this architecture. Additionally, decoder-only code generation LLMs such as CodeGen (Nijkamp et al., 2023), InCoder (Fried et al., 2023), StarCoder (Li et al., 2023a), Codex (Chen et al., 2021) and CodeLLaMa (Rozière et al., 2024) have been widely used for bug detection and repair, as well as for vulnerability repair (Wu et al., 2023a; Xia and Zhang, 2024; Yang et al., 2023b). The large context window characteristic of decoder-only models allows them to take in and utilize more context relevant to the cybersecurity task, like related vulnerabilities, reports, and code snippets.
@@
Due to the powerful natural language generation capabilities of the decoder-only architecture, researchers initially attempted to apply it to the generation of fake cyber threat intelligence (Ranade et al., 2021). Decoder-only LLMs have gained prominence in recent years, especially in 2023 and 2024 as shown in Figure 5, witnessing a surge in development and commercial adoption by leading Internet companies. For instance, Google introduced Bard (Google, 2023), while Meta unveiled LLaMa (Touvron et al., 2023). While GPT-4 and its derivative application, ChatGPT, quickly found integration into various cybersecurity tasks, these other newer models have yet to see similarly widespread adoption in the cybersecurity domain reported in the literature surveyed.